OpenStudy (anonymous):

A particle wavelength (λp) of a proton is 10-12m. What is the velocity of this proton [m/s]?

OpenStudy (anonymous):

3.96148749852*10^5

OpenStudy (anonymous):

How you get that !

OpenStudy (anonymous):

did you get the antimony reaction?

OpenStudy (anonymous):

Use the formula from deBroglie relation mpvp=h/λ so you use Vp=h/mpλ

OpenStudy (anonymous):

sorry that I didn't answered sooner... first day here and I am a little confused with all these notifications! Doriath is correct